# The growth of magnetic energy during the nonlinear phase of the subsonic and supersonic small-scale dynamo

### v1

@@ III.1 Kinematic dynamo growth
Figure 2: Kinematic growth rates, $\gamma_{\mathrm{exp}}$, as a function of plasma Reynolds number, Re, colored by $\mathcal{M}$, and markers indicating different $N_{\mathrm{res}}$. Each data point shows the median marginal posterior probability for $\gamma_{\mathrm{exp}}$ derived by fitting Equation. 2, for each combination of $(N_{\mathrm{res}},\mathcal{M},\mbox{Re})$. The vertical error bars show the 16th to 84th percentile range from the fits and the horizontal error bars show the same for the fluctuations measure in $\mbox{Re}=u_{0}\ell_{0}/\nu$ over time. Lines show fits to the median trend of $\gamma_{\mathrm{exp}}t_{0}$ versus Re for subsonic, $\gamma_{\mathrm{exp}}\propto\mbox{Re}^{1/2}$ (solid) and supersonic, $\gamma_{\mathrm{exp}}\propto\mbox{Re}^{1/3}$ (dashed) simulations, matching predictions from [46, 19], indicating the viscous scale is the SSD engine in the kinematic regime.
@@
While the scaling of $\gamma_{\mathrm{exp}}$ with Re in our simulations is in agreement with theoretical expectations, the proportionality constants we measure are systematically smaller than expected, $\sim 1/100$. A simple least-squares fit to the median values show in Fig. 2 gives $(3.22\pm 0.06)\times 10^{-2}\mbox{Re}^{1/2}$ for our $\mathcal{M}\lesssim 1$ simulations, and $(4.48\pm 0.06)\times 10^{-2}\mbox{Re}^{1/3}$ for our $\mathcal{M}>1$ simulations, indicating that the kinematic dynamo only uses $1/100$ of the hydrodynamical energy flux at the viscous scale, which we find in the next section is the same for the nonlinear dynamo. These values are significantly smaller than predicted for $\mbox{Pm}\gg 1$ flows – $37/36$ for Kolmogorov flows, and $11/60$ for Burgers [46, 19] – and instead align more closely with predictions for $\mbox{Pm}\ll 1$ flows, where $0.03$ is expected for Kolmogorov, and $5\times 10^{-3}$ for Burgers [46, 19].
@@ III.2 Nonlinear dynamo growth
Figure 3: Secular dynamo growth exponent $p_{\mathrm{nl}}$, $E_{\mathrm{mag}}\propto t^{p_{\mathrm{nl}}}$ as a function of $\mathcal{M}$ and Re. For $\mathcal{M}\lesssim 1$ simulations we unanimously recover linear growth ($p_{\mathrm{nl}}=1$; purple), and quadratic growth ($p_{\mathrm{nl}}=2$; green) for all $\mathcal{M}>1$ simulations.
Figure 4: Nonlinear growth coefficient $\alpha_{\mathrm{nl}}$ (top panel) and nonlinear phase duration normalized by $t_{0}$ (bottom panel) as a function of $\mathcal{M}$, and colored by Re. For $\mathcal{M}\lesssim 1$ we find $\alpha_{\mathrm{nl}}\propto\mathcal{M}^{3}$ (black solid line), consistent with turbulent energy-flux–regulated models of nonlinear growth: $0.05u_{0}^{3}/\ell_{0}$ from [30] (dot-dashed orange) and $(3/38)u_{0}^{3}/\ell_{0}$ from [34] (dotted orange). By contrast for $\mathcal{M}>1$, we find $\alpha_{\mathrm{nl}}\propto\mathcal{M}^{2}$ (black dashed), shallower than the flux-regulated prediction. Critically, $\alpha_{\mathrm{nl}}$ in either regime becomes independent of Re, and the cascade itself becomes the dynamo engine, with inefficient $\approx 1/100\varepsilon$ transfer that fuels the dynamo. Across all simulations the nonlinear phase persists for $t_{\mathrm{sat}}-t_{\mathrm{nl}}\approx 20t_{0}$ (solid black line in lower panel) indicating a universal duration, invariant to all plasma parameters explored in this study.
@@
Fig. 3 shows our measured growth-exponent, $p_{\mathrm{nl}}$, for $E_{\mathrm{mag}}$ during the nonlinear phase, and Fig. 4 shows both the growth-efficiency coefficient, $\alpha_{\mathrm{nl}}$, in the top panel, and duration, $t_{\mathrm{sat}}-t_{\mathrm{nl}}$, of the nonlinear phase in the bottom panel. For $\mathcal{M}\lesssim 1$ simulations, we find that growth is close to linear-in-time, $p_{\mathrm{nl}}\approx 1$, and a least-squares fit to the median results shows that the growth coefficient is $\alpha_{\mathrm{nl}}=(9.4\pm 0.7)\times 10^{-3}\mathcal{M}^{3}\approx 5\times 10^{-3}\varepsilon$, where $\varepsilon\sim u_{0}^{3}/\ell_{0}=2\mathcal{M}^{3}$ is the hydrodynamic energy-density flux rate. Critically, the growth rate is independent of Re. This is qualitatively consistent with previous works that have shown only a small, fixed fraction of $\varepsilon$ is converted into magnetic energy [31, 30], but it has never been measured with this accuracy and precision, across a broad range of plasma parameters. By comparison, Beresnyak [30] measured $\alpha_{\mathrm{nl}}/\varepsilon\approx 0.05$ (compared to our $9.4\pm 0.7\times 10^{-3}\sim 10^{-2}$) from ensemble averaged simulations, and Xu and Lazarian [34] predicted $\alpha_{\mathrm{nl}}/\varepsilon=3/38$. We show both of these predictions in Fig. 4, and while our measured efficiency is a factor of $\approx 10$ lower, as in the kinematic phase, the difference may be explained by our finite $\mbox{Pm}=1$ simulations. Thus, the overall results are consistent with the phenomenology of inefficient conversion of turbulent to magnetic energy in the nonlinear phase.
All of our $\mathcal{M}>1$ simulations yield $p_{\mathrm{nl}}\approx 2$, consistent with quadratic-in-time growth  [^1] A complementary model-comparison analysis (see End Matter) confirms that subsonic results are best described by linear growth, while quadratic growth captures the supersonic regime. of $E_{\mathrm{mag}}$, aligned with the model of Schleicher et al. [37], which predicts that $E_{\mathrm{mag}}(t)\sim E_{\mathrm{kin}}^{1/2}(t/t_{0})^{2}$ for Burgers-like turbulence. Indeed, as predicted by Schleicher et al. [37], this breaks the $p_{\mathrm{nl}}$ of the universality of the nonlinear SSD dynamo across the different $\mathcal{M}$ regimes.
@@
Furthermore, unlike the subsonic case, $\varepsilon$ no longer scales with $\mathcal{M}^{3}$, but instead follows a shallower trend $\alpha_{\mathrm{nl}}\propto\varepsilon\propto\mathcal{M}^{2}$. Empirically, the trends we measure can be understood if the conversion of $E_{\mathrm{kin}}$ into $E_{\mathrm{mag}}$ is modified by the presence of acoustic modes and shocks, i.e., if the flux is transferred on $t_{\mathrm{sc}}$, due to a fraction of the energy being directly deposited into shock heating, $\varepsilon\sim u^{2}/t_{\mathrm{sc}}\sim u^{2}c_{s}/\ell\sim\mathcal{M}^{2}$, which acts to reduce the available hydrodynamical flux by a factor of $\mathcal{M}^{-1}$. This is consistent with the idea that in the compressible regime, some fraction of the $E_{\mathrm{kin}}$ fills compressible mode degrees of freedom, which do not contribute to irreversible field amplification [28, 29, 48, 49, 25].
@@
Despite the differences in scaling and $\mathcal{M}$ dependence between the $\mathcal{M}\lesssim 1$ and $\mathcal{M}>1$ regimes, we find two aspects of the nonlinear SSD that are universal across plasma parameters and $\mathcal{M}$ regimes: firstly, the proportionality coefficient, which is associated with the nonlinear SSD efficiency, is universal across all $\mathcal{M}$ and Re, with $1/100\varepsilon$ efficiency consistent with what was found previously [30], even though $p_{\mathrm{nl}}$ is not always unity. And secondly, across the full parameter space, we find that the nonlinear phase persists for $t_{\mathrm{sat}}-t_{\mathrm{nl}}\approx(20\pm 1)\,t_{0}$, independent of $\mathcal{M}$ or Re (see Fig. 4, bottom panel), making the duration in the nonlinear dynamo also universal. This result suggests that once the nonlinear phase begins, $E_{\mathrm{mag}}\approx E_{\mathrm{kin}}(\ell_{\nu})$, the system saturates on universal time, $\approx 20t_{0}$, regardless of flow compressibility. This is likely due to the fact that even though the scaling with $\mathcal{M}$ is different between the $\varepsilon$ and hence $\alpha_{\mathrm{nl}}$, the efficiency of hydrodynamical cascade flux into the nonlinear SSD are the same for all dynamo (both kinematic and nonlinear) and plasma regimes (sub- and supersonic), $\approx 1/100$ (Fig. 4, top panel). The implication is that while the algebraic order and $\mathcal{M}$ dependency of $E_{\mathrm{mag}}$ growth depends upon the turbulent regime, the path to saturation is set by a robust and universal dynamical clock.
@@ IV Discussion and Conclusions
We present a detailed statistical analysis of small-scale dynamo (SSD) growth across a range of Reynolds Re, and sonic, turbulent sonic Mach numbers $\mathcal{M}$, focusing on the poorly-explored nonlinear phase of field growth. In order to access this regime, we (1) simulate many statistical realizations of each set of plasma parameters, allowing our measurements to become less sensitive to large statistical fluctuations that make this phase challenging to explore, and (2) use a hierarchical Bayesian model fitting technique that allows us to model the dynamo phase transitions with precision. We find that in this regime the growth rate depends on $\mathcal{M}$, but is completely liberated from any visco-resistive dynamics. $\mathcal{M}<1$ nonlinear SSDs grow $\propto t$ with an efficiency of order $\approx 1/100$ of the hydrodynamic energy flux rate from the turbulence cascade, consistent with the phenomenological models with the incompressible nonlinear SSD [16, 34]. Further, $\mathcal{M}>1$ dynamos grow $\propto t^{2}$, predicted by existing theories [19, 37], and with the same SSD efficiency $\approx 1/100$, making both the sub- and supersonic nonlinear dynamos universally inefficient in their parasitism of the hydrodynamical cascade [30].
@@
Following from the universal efficiency, we find that the nonlinear phase has an approximately universal duration of $\approx 20t_{0}$, where $t_{0}$ is the outer-scale turbulence turnover time, independent of both the Re and $\mathcal{M}$. Consequently, the time it takes systems to evolve through the nonlinear phase and reach $E_{\mathrm{mag}}$ saturation is always roughly an order of magnitude longer than the time it takes systems to reach mechanical equilibrium $\sim t_{0}$, but no more. This implies that most observed astrophysical systems, which persist for many $t_{0}$, are likely to have reached SSD saturation; only those that have undergone large perturbations very recently or that are very dynamically young are likely to be in the nonlinear phase. Why the efficiency of the nonlinear SSD is $\sim 1/100$, and the characteristic duration of the phase $\sim 20t_{0}$, rather than any other values, remains an open theoretical question.
@@
We note that our study has focused on $\mbox{Pm}=1$ plasmas, whilst many hot astrophysical plasmas are $\mbox{Pm}\gg 1$ ($\mbox{Pm}\propto T^{4}/n_{e}$, where $T$ is the plasma temperature and $n_{e}$ is the electron number density). In this regime, there may emerge a second nonlinear, secular dynamo, as the magnetic field transverses the sub-viscous range of scales [16]. The nature of dynamos in this regime, and their duration, also remain open questions for future work. Finally, our $\mathcal{M}>1$ simulations do not resolve the $u_{\ell_{s}}=c_{s}$ sonic scale, marking the transition between the sub-and-supersonic cascades [36, 50]. It may be that the nonlinear SSD transitions smoothly between a $E_{\mathrm{mag}}\sim\mathcal{M}^{3}t$ and $E_{\mathrm{mag}}\sim\mathcal{M}^{2}t^{2}$ growth phase, as in Fig. 4, which may be relevant for $\mathcal{M}\gg 1$, high-Re astrophysical plasmas. We leave the study of this regime for future, extremely high-resolution nonlinear SSD investigations.
@@ V.1 Simulation Setup
$+\left(c_{\mathrm{s}}^{2}\rho+\frac{b^{2}}{8\pi}\right)\tensor{I}-2\nu\rho\tensor{S}\bigg{]}$ $=\rho\bm{f},$ (5)
$\frac{\partial\bm{b}}{\partial t}+\bm{\nabla}\cdot\left(\bm{u}\otimes\bm{b}-\bm{b}\otimes\bm{u}\right)$ $=\eta\nabla^{2}\bm{b},$ (6)
@@
where $\rho$ is gas density, $\bm{u}$ is the gas velocity, $\bm{b}$ is the the magnetic field, $c_{\mathrm{s}}$ is sound speed, $\nu$ is the kinematic viscosity, $\eta$ is the magnetic resistivity, $\tensor{I}=\delta^{i}_{j}$ is the identity tensor and $\bm{f}$ is the turbgen forcing function [42]. We model our viscosity via the traceless strain rate tensor,
$\tensor{S}=\frac{1}{2}\left(\bm{\nabla}\otimes\bm{u}+\left(\bm{\nabla}\otimes\bm{u}\right)^{T}\right)-\frac{1}{3}\left(\bm{\nabla}\cdot\bm{u}\right)\tensor{I},$ (8)
where $\otimes$ is the tensor product $\bm{\nabla}\otimes\bm{u}\equiv\partial^{i}u_{j}$. Our simulations span a broad range of $\mathcal{M}$ and Re, and are run at a range of resolutions; see Section II.1 for details and Table 1 for a full list. $\bm{f}$ is Gaussian-random and the phases evolve in time following an Ornstein–Uhlenbeck process with correlation time set to outer-scale of the turbulence, $t_{0}$. The injection peaks on scales $\ell_{0}=L/2$, and the forcing amplitude is tuned so that the velocity field on $\ell_{0}$ stays within 5% of a chosen target value during the kinematic phase; we explore $\mathcal{M}\in[5\times 10^{-2},5]$.
@@
0.5 1500 $1.7\times 10^{-4}$ 9 $\times$ 576 $1.3_{-0.1}^{+0.1}$ $\left(1.0_{-0.2}^{+0.2}\right)\times 10^{-3}$ $1.1_{-0.1}^{+0.5}$ $14.9_{-4.5}^{+5.4}$

### v2

@@ III.1 Kinematic dynamo growth
Figure 2: Kinematic growth rates, $\gamma_{\mathrm{exp}}$, as a function of plasma Reynolds number, Re, with markers indicating different $N_{\mathrm{res}}$, and colors showing $\mathcal{M}$. Each data point shows the median marginal posterior probability for $\gamma_{\mathrm{exp}}$ derived by fitting Eqn. 2 to each plasma combination, $(N_{\mathrm{res}},\mathcal{M},\mbox{Re})$. The vertical error bars show the 16th to 84th percentile range from the fits and the horizontal error bars show the same for the fluctuations measure in $\mbox{Re}=u_{0}\ell_{0}/\nu$ over time. Lines show fits of $\gamma_{\mathrm{exp}}\propto\mbox{Re}^{1/2}$ to subsonic points (solid line) and $\gamma_{\mathrm{exp}}\propto\mbox{Re}^{1/3}$ to supersonic points (dashed line), which match predictions from [42, 43], indicating that the viscous scale is the engine for growth in the kinematic regime.
@@
While the scaling of $\gamma_{\mathrm{exp}}$ with Re in our simulations is in agreement with theoretical expectations, the proportionality constants we measure, of order $\sim 1/100$, are systematically smaller than expected. A simple least-squares fit of the dimensionless kinematic growth rate, $\gamma_{\mathrm{exp}}t_{0}$, as a function of Re, gives $(3.22\pm 0.06)\times 10^{-2}\,\mbox{Re}^{1/2}$ for our $\mathcal{M}\lesssim 1$ simulations, and $(4.48\pm 0.06)\times 10^{-2}\,\mbox{Re}^{1/3}$ for our $\mathcal{M}>1$ simulations. These fits indicate that the kinematic dynamo only converts $\sim 1/100$ of the hydrodynamical energy flux, $\varepsilon$, into magnetic energy. We show in the next section that this same conversion efficiency also holds in the nonlinear dynamo phase. These values are significantly smaller than predicted for $\mbox{Pm}\gg 1$ flows ($37/36$ for Kolmogorov flows, and $11/60$ for Burgers [42, 43]) and instead align more closely with predictions for $\mbox{Pm}\ll 1$ flows, where $0.03$ is expected for Kolmogorov, and $5\times 10^{-3}$ for Burgers [42, 43, 8].
@@ III.2 Nonlinear dynamo growth
Figure 3: Secular dynamo growth exponent $p_{\mathrm{nl}}$, measured during the nonlinear phase, as a function of $\mathcal{M}$ and Re for each plasma configuration. For $\mathcal{M}\lesssim 1$ simulations we unanimously recover linear growth ($p_{\mathrm{nl}}\approx 1$; purple), and quadratic growth ($p_{\mathrm{nl}}\approx 2$; green) for all $\mathcal{M}>1$ simulations.
[…]
@@
Fig. 3 shows our measured growth-exponent, $p_{\mathrm{nl}}$, for $E_{\mathrm{mag}}$ during the nonlinear phase, and Fig. 4 shows both the growth-efficiency coefficient, $\alpha_{\mathrm{nl}}$, in the top panel, and duration, $t_{\mathrm{sat}}-t_{\mathrm{nl}}$, of the nonlinear phase in the bottom panel. For our $\mathcal{M}\lesssim 1$ simulations, we find that growth is close to linear-in-time, $p_{\mathrm{nl}}\approx 1$, and a least-squares fit to the median results shows that the growth coefficient is $\alpha_{\mathrm{nl}}=(9.4\pm 0.7)\times 10^{-3}\mathcal{M}^{3}\approx 5\times 10^{-3}\varepsilon$, where $\varepsilon\sim u_{0}^{3}/\ell_{0}=2\mathcal{M}^{3}$ is the hydrodynamic energy flux in our simulations. Critically, the growth rate is independent of Re, suggesting that, even though we have explored only a finite range $\mbox{Re}\in[10^{3},5\times 10^{3}]$, the results are likely to be valid for all turbulent flows (where $\mbox{Re}>\mbox{Re}_{\mathrm{crit}}\approx 100$; see Kriel et al. [26]). This is qualitatively consistent with previous works that have shown only a small, fixed fraction of $\varepsilon$ is converted into magnetic energy [12, 4], but it has never been measured with this accuracy and precision, across a broad range of plasma parameters. By comparison, Beresnyak [4] measured $\alpha_{\mathrm{nl}}/\varepsilon\approx 0.05$ (compared to our $9.4\pm 0.7\times 10^{-3}\sim 10^{-2}$) from ensemble averaged $\mathcal{M}\ll 1$, $\mbox{Pm}=1$ simulations, and Xu and Lazarian [56] predicted $\alpha_{\mathrm{nl}}/\varepsilon=3/38$ for $\mathcal{M}\ll 1$, $\mbox{Pm}\gg 1$ SSDs. We show both of these predictions in Fig. 4, and while our measured efficiency is a factor of $\approx 10$ lower than both predictions, as was also the case for the growth rate measured during the kinematic phase, the difference, at least compared with Xu and Lazarian [56], can be explained by our finite $\mbox{Pm}=1$ simulations. Regardless, the overall results are consistent with the phenomenology of inefficient conversion of turbulent to magnetic energy in the nonlinear phase.
@@
All of our $\mathcal{M}>1$ simulations yield $p_{\mathrm{nl}}\approx 2$, consistent with quadratic-in-time growth  [^1] A complementary model-comparison analysis (see End Matter) confirms that subsonic results are best described by linear growth, while quadratic growth captures the supersonic regime. of $E_{\mathrm{mag}}$, aligned with the model of Schleicher et al. [41], which predicts that $E_{\mathrm{mag}}(t)\sim E_{\mathrm{kin}}^{1/2}(t/t_{0})^{2}$ for Burgers-like turbulence. Indeed, as predicted by Schleicher et al. [41], this breaks the universality of $p_{\mathrm{nl}}$ across different $\mathcal{M}$ regimes. Moreover, unlike the subsonic case, $\varepsilon$ no longer scales with $\mathcal{M}^{3}$, but instead follows a shallower trend $\alpha_{\mathrm{nl}}\propto\varepsilon\propto\mathcal{M}^{2}$. Empirically, the trends we measure can be understood if the conversion of $E_{\mathrm{kin}}$ into $E_{\mathrm{mag}}$ is modified by the presence of acoustic modes and shocks, i.e., if the flux is transferred on $t_{\mathrm{sc}}$ timescales, due to a fraction of the energy being directly deposited into shock heating, $\varepsilon\sim u^{2}/t_{\mathrm{sc}}\sim u^{2}c_{s}/\ell\sim\mathcal{M}^{2}$, which acts to reduce the available hydrodynamical flux by a factor of $\mathcal{M}^{-1}$. This is consistent with the idea that in the compressible regime, some fraction of the $E_{\mathrm{kin}}$ fills compressible mode degrees of freedom, which do not contribute to irreversible field amplification [14, 17, 47, 49, 2].
@@
Despite the differences in scaling and $\mathcal{M}$-dependence between the $\mathcal{M}\lesssim 1$ and $\mathcal{M}>1$ regimes, we find two aspects of the nonlinear SSD that are universal across plasma parameters and $\mathcal{M}$-regimes: firstly, the proportionality coefficient, which is associated with the nonlinear SSD efficiency, is universal across all $\mathcal{M}$ and Re, with $\varepsilon/100$ efficiency consistent with what was found previously [4], even though $p_{\mathrm{nl}}$ is not always unity. And secondly, across the full parameter space, we find that the nonlinear phase persists for $t_{\mathrm{sat}}-t_{\mathrm{nl}}\approx(20\pm 1)\,t_{0}$, independent of $\mathcal{M}$ or Re (see Fig. 4, bottom panel), making the duration in the nonlinear dynamo also universal. This suggests that once the nonlinear phase begins, $E_{\mathrm{mag}}\approx E_{\mathrm{kin}}(\ell_{\nu})$, the system saturates after $\approx 20t_{0}$, regardless of flow compressibility. This universal nonlinear window is visually apparent in both the bottom and inset panels of Fig. 1, where the nonlinear phase occupies a comparable interval when time is expressed as $t/t_{0}$ or $\log_{10}(t/t_{\mathrm{sc}})$. This is likely because, although $\varepsilon$ and $\alpha_{\mathrm{nl}}$ have different $\mathcal{M}$-dependencies, their ratio remains $\alpha_{\mathrm{nl}}/\varepsilon\approx 1/100$ for all dynamo (both kinematic and nonlinear) and flow regimes (sub- and supersonic) (Fig. 4, top panel). The implication is that while the algebraic order and $\mathcal{M}$ dependence of $E_{\mathrm{mag}}$ growth depends on the turbulent regime, the path to saturation is set by a robust and universal dynamical clock.
@@ IV Discussion and Conclusions
In this study we present a detailed statistical analysis of small-scale dynamo (SSD) growth across a range of hydrodynamic Reynolds, Re, and sonic Mach numbers $\mathcal{M}$, focusing on the poorly-explored nonlinear phase of field growth. In order to access this regime, we (1) simulate many statistical realizations of each set of plasma parameters, allowing our measurements to become less sensitive to large statistical fluctuations that make this phase challenging to explore, and (2) use a hierarchical Bayesian model fitting technique that allows us to precisely model the dynamo phase transitions. We find that in this regime the growth rate depends on $\mathcal{M}$, but is liberated from any visco-resistive dynamics. $\mathcal{M}<1$ nonlinear SSDs grow $\propto t$ with an efficiency of order $\approx 1/100$ of the hydrodynamic energy flux from the turbulence cascade, consistent with the phenomenological models with the incompressible nonlinear SSD [39, 56]. Further, $\mathcal{M}>1$ dynamos grow $\propto t^{2}$, predicted by existing theories [43, 41], and with the same SSD efficiency $\approx 1/100$, making both the sub- and supersonic nonlinear dynamos universally inefficient in their parasitism of the hydrodynamical cascade [4].
@@
Following from the universal efficiency, we find that the nonlinear phase has an approximately universal duration of $\approx 20t_{0}$, where $t_{0}$ is the outer-scale turbulence turnover time, independent of both the Re and $\mathcal{M}$. Consequently, the time it takes systems to evolve through the nonlinear phase and reach $E_{\mathrm{mag}}$ saturation is always roughly an order of magnitude longer than the time it takes systems to reach mechanical equilibrium $\sim t_{0}$, but no more. This implies that most observed astrophysical systems, which persist for many $t_{0}$, are likely to have reached SSD saturation; only those that have undergone large perturbations very recently, or that are very dynamically young, are likely to be in the nonlinear phase. Why the efficiency of the nonlinear SSD is $\sim 1/100$, and the characteristic duration of the phase $\approx 20t_{0}$, rather than any other values, remains an open theoretical question.
We note that our study has focused on $\mbox{Pm}=1$ plasmas, while many hot astrophysical plasmas are $\mbox{Pm}\gg 1$ ($\mbox{Pm}\propto T^{4}/n_{e}$, where $T$ is the plasma temperature and $n_{e}$ is the electron number density). Our choice of $\mbox{Pm}=1$ is motivated by the fact that in this study we target the nonlinear phase that begins once $E_{\mathrm{mag}}\approx E_{\mathrm{kin}}$ on the viscous scale, and the backreaction that operates on dynamically important (viscous and larger) scales. In $\mbox{Pm}\gg 1$ systems, theory predicts that the nonlinear phase studied here may be preceded by another nonlinear, secular dynamo phase, where the magnetic field traverses the sub-viscous range of scales before reaching the viscous scale [39]. Our results therefore provide a useful anchor for future studies that explore this earlier stage of nonlinear evolution.
@@
Finally, we note that our $\mathcal{M}>1$ simulations do not resolve a $u_{\ell_{s}}=c_{s}$ sonic scale, where the velocity spectrum transitions between a subsonic and supersonic cascade [16, 3]. It may be that the nonlinear SSD transitions smoothly between the $E_{\mathrm{mag}}\sim\mathcal{M}^{3}t$ and $E_{\mathrm{mag}}\sim\mathcal{M}^{2}t^{2}$ growth phase, as in Fig. 4, which may be relevant for $\mathcal{M}\gg 1$, $\mbox{Re}\gg 1$ astrophysical plasmas. We leave the study of this regime for future, extremely high-resolution nonlinear SSD investigations.
@@ V.1 Simulation Setup
$+\left(c_{\mathrm{s}}^{2}\rho+\frac{b^{2}}{8\pi}\right)\tensor{\bm{I}}-2\nu\rho\tensor{\bm{S}}\bigg]$ $=\rho\bm{f},$ (5)
$\frac{\partial\bm{b}}{\partial t}+\bm{\nabla}\cdot\left(\bm{u}\otimes\bm{b}-\bm{b}\otimes\bm{u}\right)$ $=\eta\bm{\nabla}^{2}\bm{b},$ (6)
@@
where $\rho$ is gas density, $\bm{u}$ is the gas velocity, $\bm{b}$ is the the magnetic field, $c_{\mathrm{s}}$ is sound speed, $\nu$ is the kinematic viscosity, $\eta$ is the magnetic resistivity, $\tensor{\bm{I}}=\delta^{i}_{j}$ is the identity tensor. We model our viscosity via the traceless strain rate tensor,
$\tensor{\bm{S}}=\frac{1}{2}\left(\bm{\nabla}\otimes\bm{u}+\left(\bm{\nabla}\otimes\bm{u}\right)^{T}\right)-\frac{1}{3}\left(\bm{\nabla}\cdot\bm{u}\right)\tensor{\bm{I}},$ (8)
where $\otimes$ is the tensor product $\bm{\nabla}\otimes\bm{u}\equiv\partial^{i}u_{j}$. Our simulations span a broad range of $\mathcal{M}$ and Re, and are run at a range of resolutions; see Section II.1 for details and Table 1 for a full list. The forcing function $\bm{f}$ is Gaussian-random and the phases are evolved in time using the turbgen implementation [15] of an Ornstein-Uhlenbeck process, where the correlation time sets to outer-scale of the turbulence, $t_{0}$. We set the injection to peak on $\ell_{0}=L/2$, and tune the forcing amplitude so that the velocity field on $\ell_{0}$ stays within 5% of a chosen target value during the kinematic phase; we explore $\mathcal{M}\in[5\times 10^{-2},5]$.
@@
0.5 1500 $1.7\times 10^{-4}$ 6 $\times$ 576 $1.3_{-0.1}^{+0.2}$ $\left(0.7_{-0.6}^{+0.2}\right)\times 10^{-3}$ $1.1_{-0.1}^{+0.5}$ $14.9_{-4.5}^{+5.4}$